Oct 23, 2014 Calorimetry

From GlueXWiki
Jump to: navigation, search

Video Conferencing Information

Meeting Time: 11:30 a.m. Eastern (until switch from DST and then return to 11:00am)

  1. To join via a Web Browser, go to the page [1] https://bluejeans.com/907185247.
  2. To join via Polycom room system go to the IP Address: (bjn.vc) and enter the meeting ID: 907185247.
  3. To join via phone, use one of the following numbers and the Conference ID: 907185247.
    • US or Canada: +1 408 740 7256 or
    • US or Canada: +1 888 240 2560
  4. Upon connection all microphones are automatically muted. To unmute your mike on a Polycom or equivalent unit, enter *4. Unmuting on a computer is trivial as there is a microphone button than can be clicked.
  5. More information on connecting to bluejeans is available.

Participant Direct Lines

  • JLab Phone: in CC F326 is 757-269-6460 (usual room)
  • JLab Phone in CC L207 is 757-269-7084
  • Phone in the Regina Video-conference Suite is 306-585-4204
  • Athens Phone: in Christina's office is 011-30-210-727-6947

Action Items

  1. Add iron shields to top cosmic-ray counters (Elton)
  2. Check into procedure for new MOUs (Elton)
  3. Check specification of voltage files used during a data set (Elton) See GlueX-doc-2551

Tentative Agenda

  1. Announcements
  2. Action Items
  3. On call cell phone
  4. Preparations for beam: Checklists, Accessing Onsite Webpages From Offsite
  5. FCAL HDFCAL log book
    1. FCAL Monitoring Plots (Matt)
  6. BCAL HDBCAL log book
    1. Spectra from Available Cosmic Runs (Andrei)
  7. Reconstruction and Simulation
    1. Dark hits in BCAL simulation (Will L.)
  8. Calibration
  9. Any other business


Attendees: Elton, Will M., Manuel, Mark (JLab); Matt (IU); Zisis, Andrei, Teigan (UofR); Christina (Athens); Will L.(CMU)

  1. Announcements
  2. Action Items
  3. FCAL HDFCAL log book
    1. Preparations for Beam (Manuel)
      • Updating shift-taker's screens
      • Few bases spontaneously set the voltage to 2200 V and cannot be addressed individually
      • It is possible to power cycle one strand (of 100 bases) to set the voltage to zero and leave it there
      • The power cycle is an expert operation and can be done remotely. Unfortunately, due to some cabling issues not all strands can be accessed
      • Documentation is being added with instructions to shift takers regarding this issue
      • Matt: Symptoms point to a software/EPICs issue, as this has not been seen before integration to EPICS
      • Mark volunteered to meet with Hovanes and Manuel to try to isolate this problem.
    2. FCAL Monitoring Plots (Matt)
      • File shows various monitoring plots for the FCAL to be included into ROOTSPY.
      • Kei is aware of the updates and will integrate them into the online system.
  4. BCAL HDBCAL log book
    1. Mark: A change to the input to the unified hit object. The time was taken from the TDC when available, otherwise from the ADC. Since the two times are not referenced to the same time, this leads to disparate time values. Mark will change the code, for the moment, to always use the Flash ADC time. It can be changed back later when the TDC time is better understood.
    2. Rates
      • Mark is currently working on the transfer of rates to EPICS for the BCAL. Considering changing the variable names, but needs to consult with Hovanes. Output of arrays for the rates is also problematic as north and south rates are located in different crates.
      • Work on rates for ST and PS are complete
      • Next will integrate the TOF rates
      • Work on the FCAL will follow, but the rates are taken from the FCAL Flashes (instead of discriminators), which is a completely different system. Also, David is worried that use of the same protocol used for the TDCs might interfere with the DAQ system. An alternative scheme of client/server could be used but Mark needs to learn how it works.
    3. Spectra from Available Cosmic Runs (Andrei)
      • Qualitative agreement with expectations from simulation
      • Concern that the threshold is too high. From the plots it appears to be at about 200 counts. The peaks at about 350 counts are fine, but others at lower values may be cut off by the threshold
      • Mark: Estimate that about 1/4 of the channels could be calibrated reliably with the current threshold.
      • Suggestions for updates: a) Check pedestals, especially whether the 20 samples for calculating the pedestal are not affected by the pulses b) check that the summing of 25-70 samples is not missing part of the pulse. c) number of entries in at least one plot is anomalous d) determine peak values from all channels to understand trends in the data.
  5. Reconstruction and Simulation
    1. Dark hits in BCAL simulation (Will L.)
      • Study of dark pulses. Dark pulses in the real data are not very apparent, but very large in the simulation.
      • Check to see if it is possible to turn the dark pulses off in simulation. This leads to problems, as the simulation is not intended to run this way
      • Simulation: dark pulse "problem" increases for the outer layers. Not noticed in the real data, but Mark will check this
      • Mark: pedestal sigma in data ~ 1.5 counts, but sigma ~ 5 counts in simulation. This could be a hint to problem
      • Mark: Should investigate whether changing the amplitude of pixels can bring the simulation closer to data.

Elton: The dark rate of the SiPMs was well known at the time the simulation was written. However, the relative gain was quite uncertain, so this is a natural place to check for a discrepancy.

  1. Calibration
  2. Any other business
    • Zisis will send Elton the schedule for Regina travel for reference.